# Spatio-Temporal Adaptive Diffusion Models for EEG Super-Resolution in Epilepsy Diagnosis

### v2

@@ 2.3.3 Multi-scale Transformer Denoising Module
EEG, limited by the number and placement of electrodes, struggles with spatial resolution. High-density EEG devices have shown promise in tasks like epileptic source localization and BCI. The multi-scale Transformer denoising module is designed to enhance the spatial resolution of existing EEG devices. Traditional diffusion models, which typically use UNet, are effective in image generation but limited in handling the long-sequence dependencies of time-series data like EEG. Inspired by [36, 37], the MTD module employs a Transformer backbone to more effectively extract spatio-temporal features. Additionally, multi-scale 1D convolution blocks are introduced to capture multi-scale temporal features in the reverse denoising process.
As shown in the blue part of Fig. 1, the MTD module mainly consists of position encoding, multi-scale 1D convolution blocks and diffusion Transformer blocks. Each diffusion Transformer block includes a normalization layer, a MSA layer, a cross-attention layer and a feed-forward network. The MTD module takes , timestep and spatio-temporal condition as inputs, outputting the predicted noise at timestep . The multi-scale 1D convolution blocks extract temporal features at different scales from the latent vectors using 1D convolution layers with various kernel sizes and concatenate them along the feature dimension. The output of this convolution path can be represented as:
where Conv denotes performing a 1D convolution on each channel sequence and BN represents batch normalization. denotes the output of different 1D convoluntion layer and denotes the number of 1D convoluntion layers. Subsequently, each row of is treated as a token and input into the diffusion Transformer block. Through the MSA layer, information from different scales is integrated, and global spatio-temporal dependencies are captured. For the input , where is the sequence length and is the feature dimension. The output of MSA layer can be expressed as
@@
where are projection matrices with learnable parameters. is the scaling factor. Finally, the output is processed by a feed-forward network to obtain the final fused feature representation, and a standard linear decoder is used to output a predicted noise. This enables the reparameterization trick to obtain the denoised latent for the previous time step. This mechanism effectively integrates the temporal and spatial features of EEG, providing rich context for subsequent denoising and reconstruction tasks, thereby skillfully achieving the mapping from LR EEG to HR EEG.
@@ 2.4 Loss Function
In the forward process of the diffusion model, we gradually add Gaussian noise to the clean input to generate a series of noisy data . In the reverse process, the traditional method is to train a denoising model to predict the state of the previous time step. Diffusion models draw on the principles of variation method [38], optimizing model parameters by maximizing a variational lower bound based on the evidence lower bound (ELBO):
where KL denotes the Kullback-Leibler divergence. However, directly predicting may lead to a complex and unstable training process. To simplify the training process, we use the denoising model to predict the noise at the current time step , and then minimize the mean squared error loss between the predicted noise and the ground-truth noise for each time step to learn the model parameters, where x denotes the input LR EEG. The final loss function can be expressed as:
where is the proposed multi-scale Transformer denoising model and represents the proposed STC module. Algorithm 1 displays the complete training procedure with the final loss function.
@@ 3.1 Dataset and Preprocessing
In this study, we utilized a publicly available dataset (Localize-MI [39]) for model training and validation. The dataset consists of high-density EEG data collected from seven drug-resistant epilepsy patients (mean age=35.1, age=5.4), comprising a total of 61 sessions (average session duration per subject=8.71, session duration per subject=2.65). Additionally, the dataset includes spatial locations of stimulus contacts in MRI space, MNI152 space [40], and Freesurfer surface space [41], corresponding to each subject, as well as the digitized positions of 256 scalp EEG electrodes. We employed the MNE software for preprocessing of the raw data. Following the work of [39], we applied high-pass filtering to the continuous EEG time series to remove low-frequency irrelevant information. To eliminate the large line noise caused by direct current, we used notch filtering at 50, 100, 150, and 200 Hz. Subsequently, we segmented the raw data into epochs of fixed duration (350 ms) based on the provided annotation events in the dataset. Moreover, the dataset was recorded using the EGI NA-400 amplifier (Electrical Geodesics) at a sampling rate of 8000 Hz for 256-channel EEG data, and deployed with the Geodesic Sensor Net using the HydroCel CleanLead electrode distribution system.
@@ 3.2.1 Implement Details
The proposed STADMs are implemented using the PyTorch framework, and the corresponding experiments were conducted on an Nvidia RTX A800 GPU. During model training, the Adam optimizer was employed with a batch size of 32, a learning rate of , and 300 training epochs. In the diffusion model framework, the forward process was set to 1000 time steps, with noise levels following sine or cosine schedules, and Gaussian noise was used for the noise type. In the multi-scale 1D convolution blocks, we employed 1D convolution layers with kernel sizes of 3, 5, 7, and 9, respectively. In the diffusion Transformer block, the number of attention heads was set to 16, and the hidden dimension was 64.
@@ 3.2.2 Metrics
To evaluate the quality of the synthetic EEG signals, we employed four quantitative metrics: Pearson Correlation Coefficient (PCC), Mean Absolute Error (MAE), Normalized Mean Squared Error (NMSE), and Signal-to-Noise Ratio (SNR) [43]. These metrics provide comprehensive insights into the accuracy and reliability of the reconstructed signals compared to the real signals.
To quantitatively evaluate the performance improvement of SR EEG in downstream tasks, we used classification performance metrics including Accuracy (ACC), Precision, Recall, and F1 Score. Additionally, to assess the performance of SR EEG in source localization, we used Localization Error [7] as a quantitative measure of performance improvement in source localization.
3.3 Performance Evaluation on EEG SR Reconstruction
In this section, we evaluate the performance of the proposed STADMs for EEG spatial super-resolution reconstruction, primarily analyzing the generated results through various quantitative metrics and scalp topography maps.
@@ 3.3.1 Comparison With Other Existing EEG SR Methods
Fig. 2 shows quantitative comparison results between different methods. As observed, the proposed STADMs achieve the best performance, significantly outperforming other existing competitive methods. The highest PCC and SNR values indicate that the reconstructed results of STADMs maintain high signal quality and correlation. Moreover, STADMs significantly outperform other methods in terms of MAE and NMSE, indicating lower errors and higher accuracy in reconstructing SR EEG. Additionally, the gaps between STADMs and other methods are substantial across all quantitative metrics. Through the experimental results, it can be concluded that compared to existing EEG SR methods, the proposed STADMs effectively reconstruct high spatial resolution(256 channels) EEG, achieving significant spatial resolution enhancement of LR EEG.
@@ 3.3.3 Qualitative results of Synthetic EEG
To observe the differences between synthetic SR EEG and LR EEG at different channel levels (16, 32, 64, and 128) more intuitively, we analyze the scalp topomaps of EEG in different frequency bands.We first use the Multitaper method to extract the power spectral density (PSD) of LR EEG and corresponding synthetic SR EEG in five frequency bands, including (0.5-4 Hz), (4-8 Hz), (8-13 Hz), (13-30 Hz), and (30-40 Hz or higher) [45]. Finally, we use the MNE tool to obtain scalp topographies in different frequency bands. Fig. 4 compares scalp topographies across different frequency bands between EEG at various channel levels and model-generated 256-channel SR EEG. Each column represents a frequency band, and each row represents a type of EEG. As spatial resolution increases, the active regions on the scalp topographies become smaller and clearer. For instance, the 16-channel LR EEG shows large, blurry active regions in both low (, , ) and high-frequency bands (, ), whereas the 256-channel SR EEG accurately localizes active sources in high-frequency bands. The synthesized SR EEG retains the frequency characteristics of the original signal across all bands, effectively reconstructing their spatial distribution. This demonstrates the strong generalization of the model across different frequency bands and its ability to stably reconstruct EEG signals. Overall, the proposed method generates SR EEG with richer spatial detail, more accurately describing and locating active regions on the scalp, whereas LR EEG results in coarse active regions, making it challenging to pinpoint active brain areas.
@@ 3.4 Classification Evaluation on the Synthetic SR EEG
| freq bands | EEG | Accuracy (AVGSTD %) | ||
| scaling factor 4 | scaling factor 8 | scaling factor 16 | ||
@@
We also analyzed classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to n 5 features (where n is the number of channels), and classification is based on these features. Table 2 shows classification results for LR and SR EEG across different frequency bands and scaling factors. SR EEG outperforms LR EEG in all frequency bands. As the scaling factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands (, , ) have worse performance than high-frequency bands (, ), but the average accuracy improvement is higher in low-frequency bands (5.06%) compared to high-frequency bands (4.95%). This indicates that the proposed method better preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
@@ 3.5 Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [47]. To evaluate the effectiveness of the proposed method in enhancing the spatial resolution of EEG and improving downstream task performance, we assess it using the source localization task, incorporating both quantitative and qualitative analyses. For quantitative analysis, we use localization error [7] as the metric, quantified as the Euclidean distance between the source localization results and the true source positions, to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [48], an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [49]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
@@
Fig. 7 shows the visualization results of different EEG data types on source localization tasks for some subjects. In these visualizations, green circles represent the stimulus source location ground-truth, red circles indicate the original HR EEG results, yellow circles denote the synthesized SR EEG results, and blue circles show the results from the LR EEG. The LR EEG results exhibit significant deviations from the ground truth, whereas SR EEG results show minimal deviation, closely matching HR EEG performance. This demonstrates that the proposed method effectively enhances spatial resolution and signal quality, significantly improving source localization accuracy for LR EEG. Overall, the method not only enhances EEG spatial resolution but also supports high-precision EEG imaging analysis in clinical and research applications.
@@ 4 Discussion and Conclusion
Experimental results show that while the proposed STADMs outperform existing methods in reconstruction performance, larger scaling factors reveal a significant difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 3. We identified several possible reasons for this: i) At larger scaling factors, the available spatial information for the model is limited, making it difficult to accurately capture the spatial relationships between channels; ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model; iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of other physiological signals. In the future, we will further optimize this framework and apply it to practical clinical scenarios in epilepsy diagnosis, aiding doctors in the preoperative assessment for epilepsy focus resection surgery.

### v3

@@ 2.3.3 Multi-scale Transformer Denoising Module
EEG, limited by the number and placement of electrodes, struggles with spatial resolution. High-density EEG devices have shown promise in tasks like epileptic source localization and BCI. The multi-scale Transformer denoising module is designed to enhance the spatial resolution of existing EEG devices. Traditional diffusion models, which typically use UNet, are effective in image generation but limited in handling the long-sequence dependencies of time-series data like EEG. Inspired by [44, 45, 46], the MTDM employs a Transformer backbone to more effectively extract spatio-temporal features. Additionally, multi-scale 1D convolution blocks [47] are introduced to capture multi-scale temporal features in the reverse denoising process .
As shown in the blue part of Fig. 1, the MTDM mainly consists of position encoding, multi-scale 1D convolution blocks and diffusion Transformer blocks. Each diffusion Transformer block includes a normalization layer, a MSA layer, a cross-attention layer and a feed-forward network. The MTDM takes , timestep and spatio-temporal condition as inputs, outputting the predicted noise at timestep . The multi-scale 1D convolution blocks extract temporal features at different scales from the latent vectors using 1D convolution layers with various kernel sizes and concatenate them along the feature dimension. This convolution path yields an output that can be represented as:
where Conv denotes performing a 1D convolution on each channel sequence and BN represents batch normalization. denotes the output of different 1D convoluntion layer and denotes the number of 1D convoluntion layers. Subsequently, each row of is treated as a token and input into the diffusion Transformer block. Through the MSA layer, information from different scales is integrated, and global spatio-temporal dependencies are captured. For the input , where is the feature dimension and denotes the sequence length. The output of MSA layer can be expressed as
@@
where are projection matrices with learnable parameters. is the scaling factor. Finally, the fused feature representation is obtained by processing the attention results through a feed-forward network, and a standard linear decoder is used to output a predicted noise. This enables the reparameterization trick to obtain the denoised latent for the previous time step. This mechanism effectively integrates the temporal and spatial features of EEG, providing rich context for subsequent denoising tasks, thereby skillfully achieving the adaptive generation of SR EEG.
@@ 2.4 Loss Function
In the forward diffusion process, diffusion models progressively add Gaussian noise to the clean input to generate a series of noisy data . In the reverse process, the traditional method is to train a denoising model to predict the state of the previous time step. Diffusion models draw on the principles of variation method [48], optimizing model parameters by maximizing a variational lower bound based on the evidence lower bound (ELBO):
where KL denotes the Kullback-Leibler divergence. However, directly predicting may lead to a complex and unstable training process. To simplify the training process, the MTDM is employed to predict the noise at each time step , and the MSE loss is minimized to learn the model parameters between the predicted noise and actual noise , where denotes the input LR EEG. The final loss function can be expressed as:
where is the proposed multi-scale Transformer denoising model and represents the proposed STCM. Algorithm 1 presents the entire training procedure along with the final loss function.
@@ 3.1 Dataset and Preprocessing
In this work, we leveraged a publicly accessible dataset (Localize-MI [49]) for model training and validation. The dataset consists of high-density EEG data collected from seven drug-resistant epilepsy patients (mean age=35.1, age=5.4), comprising a total of 61 sessions (average session duration per subject=8.71, session duration per subject=2.65). Additionally, the dataset contains the spatial locations of stimulus contacts in MRI space, Freesurfer surface space [50], and MNI152 space [51]for each subject, along with the digital coordinates of 256 scalp EEG electrodes. We employed the MNE software for preprocessing of the raw data. Following the work of [49], we applied high-pass filtering to the continuous EEG time series to remove low-frequency irrelevant information. To eliminate the large line noise caused by direct current, we used notch filtering at 50, 100, 150, and 200 Hz. Subsequently, we segmented the raw data into epochs of fixed duration (350 ms) based on the provided annotation events in the dataset. Moreover, the dataset was recorded using the EGI NA-400 amplifier (Electrical Geodesics) at a sampling rate of 8000 Hz for 256-channel EEG data, and deployed with the Geodesic Sensor Net using the HydroCel CleanLead electrode distribution system.
@@ 3.2.1 Implement Details
The proposed STADMs are implemented using the PyTorch framework, and the corresponding experiments were conducted on an Nvidia RTX A800 GPU. With the Adam optimizer, training of the model was carried out using a learning rate of , 300 epochs, and a batch size of 32. In the diffusion model framework, the forward process was set to 1000 time steps, with noise levels following sine or cosine schedules, and Gaussian noise was used for the noise type. In the multi-scale 1D convolution blocks, we employed 1D convolution layers with respective kernel sizes of 3, 5, 7, and 9. In the diffusion Transformer block, the number of attention heads was set to 16, and the hidden dimension was 64.
@@ 3.2.2 Metrics
To evaluate the quality of the synthetic EEG signals, we employed four quantitative metrics: Pearson Correlation Coefficient (PCC), Normalized Mean Squared Error (NMSE), Signal-to-Noise Ratio (SNR) [53], and Mean Absolute Error (MAE). These metrics offer a thorough understanding of the accuracy and reliability of the synthetic signals in comparison to the real signals.
The performance improvement of SR EEG in downstream tasks was quantitatively assessed using classification performance metrics including Accuracy (ACC), Precision, F1 Score, and Recall. Additionally, we used Localization Error [7] as a quantitative measure of performance improvement in source localization.
3.3 Evaluation of EEG SR Reconstruction Performance
In this section, the performance of the proposed STADMs for EEG spatial super-resolution reconstruction is evaluated through various quantitative metrics and scalp topography maps.
@@ 3.3.1 Comparison With Other Existing EEG SR Methods
Fig. 2 shows quantitative comparison results between different methods. As indicated, the proposed STADMs exhibit superior performance, significantly surpassing other existing competitive methods. The highest PCC and SNR values indicate that the reconstructed results of STADMs maintain high signal quality and correlation. Moreover, STADMs significantly outperform other methods in terms of MAE and NMSE, indicating lower errors and higher accuracy in reconstructing SR EEG. Additionally, the gaps between STADMs and other methods are substantial across all quantitative metrics. Through the experimental results, it can be concluded that compared to existing EEG SR methods, the proposed STADMs effectively reconstruct high spatial resolution(256 channels) EEG, achieving significant spatial resolution enhancement of LR EEG.
@@ 3.3.3 Qualitative results of Synthetic EEG
To observe the differences between synthetic SR EEG and LR EEG at different channel levels (16, 32, 64, and 128) more intuitively, we analyze the scalp topomaps of EEG in different frequency bands. Initially, the Multitaper method is employed to extract the power spectral density (PSD) of EEG with different spatial resolutions across five frequency bands, including (30-40 Hz or higher), (13-30 Hz), (8-13 Hz), (4-8 Hz), and (0.5-4 Hz) [55]. Finally, we use the MNE tool to obtain scalp topographies in different frequency bands. Fig. 4 compares scalp topographies across different frequency bands between EEG at various channel levels and model-generated 256-channel SR EEG. Each column represents a frequency band, and each row represents a type of EEG. As spatial resolution increases, the active regions on the scalp topographies become smaller and clearer. For instance, the 16-channel LR EEG shows large, blurry active regions in both low (, , ) and high-frequency bands (, ), whereas the 256-channel SR EEG accurately localizes active sources in high-frequency bands. The synthesized SR EEG retains the frequency characteristics of the original signal across all bands, effectively reconstructing their spatial distribution. This demonstrates the strong generalization of the model across different frequency bands and its ability to stably reconstruct EEG signals. Overall, the proposed method generates SR EEG with richer spatial detail, more accurately describing and locating active regions on the scalp, whereas LR EEG results in coarse active regions, making it challenging to pinpoint active brain areas.
@@ 3.4 Classification Evaluation on the Synthetic SR EEG
| freq bands | EEG | scaling factor / Accuracy (AVGSTD %) | ||
| 4 | 8 | 16 | ||
@@
We also analyzed classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to features (where c denotes the number of channels), and classification is performed based on these features. Table 2 shows classification results for LR and SR EEG across different frequency bands and scaling factors. SR EEG outperforms LR EEG in all frequency bands. As the scaling factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands (, , ) have worse performance than high-frequency bands (, ), but the average accuracy improvement in low-frequency bands (5.06%) surpasses that in high-frequency bands (4.95%). This suggests that the proposed STADMs more Adequately preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
@@ 3.5 Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [57]. To determine the impact of the proposed STADMs in enhancing EEG spatial resolution and its effectiveness in improving downstream task performance, we assess it using the source localization task, incorporating both quantitative and qualitative analyses. We utilize localization error [7] for quantitative analysis, defining it as the Euclidean distance between the source localization results and the ground-truth positions. This metric is employed to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [58], an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [59]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
@@
Fig. 7 shows the visualization results of different EEG data types on source localization tasks for some subjects. In these visualizations, green circles represent the stimulus source location ground-truth, red circles indicate the original HR EEG results, yellow circles denote the synthesized SR EEG results, and blue circles show the results from the LR EEG. The LR EEG results display considerable deviations from the ground truth, while the SR EEG results show minimal deviation, closely aligning with HR EEG performance. This demonstrates that the proposed method effectively enhances spatial resolution and signal quality, significantly improving source localization accuracy for LR EEG. Overall, the method not only enhances EEG spatial resolution but also supports high-precision EEG imaging analysis in clinical and research applications.
@@ 4 Discussion and Conclusion
Experimental results show that while the proposed STADMs outperform existing methods in reconstruction performance, larger scaling factors reveal a significant difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 3. We identified several possible reasons for this: i) With larger scaling factors, the available spatial information for the model is constrained, making it challenging to accurately capture the spatial relationships between channels. ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model. iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of other physiological signals. In the future, we will further optimize this framework and apply it to practical clinical scenarios in epilepsy diagnosis, aiding doctors in the preoperative assessment for epilepsy focus resection surgery.